A turbine operating under steady flow conditions receives steam at the


following state : pressure 13.8 bar ; specific volume 0.143 m3/kg ;


internal energy 2590 kJ/kg ; velocity 30 m/s. The state of the steam


leaving the turbine is : pressure 0.35 bar ; specific volume 4.37 m3/kg


; internal energy 2360


kJ/kg ; velocity 90 m/s. Heat is lost to the surroundings at the rate of


0.25 kJ/s. If the rate of steam flow is 0.38 kg/s, what is the power


developed by the turbine ?

1
Expert's answer
2022-02-16T19:23:02-0500

Here,


u - specific internal energy at inlet and exit

P- Pressure at inlet and exit.

V - specific volume at inlet and e

M- mass flow rate through turbine

C - velocity at inlet and exit

Q - Rate of heat transfer between turbine and surroundings

W - rate of work transfer


inlet power

"=m(u+pv+\frac{c_2}{^2})-Q"

"=0.38(2\n.590\u00d710^6+13.8\u00d710^5\u00d70.143+450)-250"


"=1059110"


Outlet power

"0.38(2.36\u00d710^6+0.35\u00d710^5\u00d74.37+450)+w"


"=956460+w"


"1059110=956460+w"

"w=1059110-956460"

"w=102650" W